# Segment Any Text: A Universal Approach for Robust, Efficient and Adaptable Sentence Segmentation

### v1

@@ Abstract
Segmenting text into sentences plays an early and crucial role in many NLP systems. This is commonly achieved by using rule-based or statistical methods relying on lexical features such as punctuation. Although some recent works no longer exclusively rely on punctuation, we find that no prior method achieves all of (i) robustness to missing punctuation, (ii) effective adaptability to new domains, and (iii) high efficiency. We introduce a new model — Segment any Text (SaT) — to solve this problem. To enhance robustness, we propose a new pretraining scheme that ensures less reliance on punctuation. To address adaptability, we introduce an extra stage of parameter-efficient fine-tuning, establishing state-of-the-art performance in distinct domains such as verses from lyrics and legal documents. Along the way, we introduce architectural modifications that result in a threefold gain in speed over the previous state of the art and solve spurious reliance on context far in the future. Finally, we introduce a variant of our model with fine-tuning on a diverse, multilingual mixture of sentence-segmented data, acting as a drop-in replacement and enhancement for existing segmentation tools. Overall, our contributions provide a universal approach for segmenting any text. Our method outperforms all baselines — including strong LLMs — across 8 corpora spanning diverse domains and languages, especially in practically relevant situations where text is poorly formatted. [^1] Our models and code, including documentation, are available at https://huggingface.co/segment-any-text under the MIT license.
@@ 2 Introduction
However, the performance of WtP as the current state-of-the-art (SoTA) model and all other segmenters is still poor on texts from more challenging domains. This includes, among others, user-generated text such as tweets and highly heterogeneous domains such as lyrics. Segmenting these texts is challenging because of missing and/or extra punctuation, inconsistent spacing, and especially irregular casing. Furthermore, nearly all existing systems, including WtP, require the specification of the texts’ language at inference time. This necessitates an additional preprocessing step of language identification, which often proves to be imperfect, particularly with user-generated content (Lui and Baldwin, 2014; Sterner and Teufel, 2023). Moreover, this necessity limits their applicability to code-switched text, as assigning a single language code to mixed-language inputs is impossible.
@@
Figure 1 shows that the standard 3-layer version of SaT outperforms the current state-of-the-art, WtP, while achieving a $\approx 3$x reduction in inference time. Overall, we present several innovations that overcome each of the shortcomings of previous methods, culminating in a universal model for sentence segmentation. We provide some examples of our model’s predictions in Figure 2.
@@
1) We introduce Segment any Text (SaT), an efficient method for sentence segmentation that can reliably segment text across 85 languages regardless of lexical features such as punctuation or casing. 2) We show how our models can be adapted to different domains via data-efficient means, requiring only a minimal set (e.g., 16) of sentence-segmented examples. 3) We train and release SaT models in five sizes, covering 85 languages, and demonstrate state-of-the-art performance across 8 corpora, even outperforming newly introduced strong LLM baselines.
@@ 3.2 Where’s the Point (WtP)
Robustness to corruptions. To make WtP less reliant on punctuation, Minixhofer et al. (2023) randomly remove some punctuation during training. In addition, they predict the likelihood of commonly occurring punctuation as an auxiliary objective. While this helps make WtP models less reliant on punctuation, we still find that WtP models have major issues when text is inconsistently formatted, especially irregular casing.
@@ 4 SaT: Segment any Text
Robustness to corruptions. We randomly remove common punctuation-only tokens with probability $p$ and use the auxiliary punctuation-prediction objective during training. Further, we randomly remove all casing and punctuation in 10% of samples within a batch during training. The resulting model, Segment any Text (SaT), already shows strong segmentation performance at improved efficiency.
Still, to further improve SaT, we continue training it on a Supervised Mixture of already-segmented sentences. [^2] Although more domain-specific, this is motivated by the strong multi-task learning abilities of T5 (Raffel et al., 2019). To be even less dependent on patterns such as punctuation, spaces, and casing, we augment the data by introducing several additional corruption schemes, resulting in our more specialized model, SaT+SM.
Our first corruption scheme removes all casing and punctuation tokens for all text. Secondly, we add randomness to corruption in as many situations as we deem useful, e.g., user-generated text in tweets or forums. This includes duplicating punctuation, removing punctuation, lowercasing, and removing/adding spaces between sentences. Finally, we also use clean, non-corrupted text. We then sample uniformly across these three categories. For details, see Appendix A.2.
@@
Code- switching C.f. Table 10. Reddit posts for German-English (de-en); data for 3 additional language pairs is taken by concatenating code-switching sentences from bilingual transcriptions. We treat all data as transcriptions, removing all punctuation and casing; we only keep sentences with at least one token of each language. Deuchar (2009), Osmelak and Wintner (2023), Nguyen and Bryant (2020) and Çetinoğlu (2017)
@@ 6.1 Performance on Clean Text
Table 2 shows evaluation results on clean text, averaged over OPUS100, UD, and Ersatz on a diverse selection of languages, including an average over 81 languages.777 For the average, we only consider languages with datasets with both train and test sets for a fair comparison. While we evaluate on 85 languages, this is the case in 81 languages. We categorize methods into (i) multilingual, which take only text as input, and (ii) monolingual, which additionally rely on a language code or, in the case of WtPT, WtPPUNCT, and SaT+LoRA, are adapted to a target domain.
@@ 6.3 Performance on Challenging Domains
Code-switching. The results in Table 5 reveal that WtP achieves an average F1 score of $29.1\%$, while the highest-performing LLM scores $43.4\%$. SaT and SaT+SM achieve average F1 of $48.2\%$ and $54.4\%$, respectively. SaT+LoRA continues to improve performance, achieving $61.7\%$. To the best of our knowledge, this is the first evaluation of sentence segmentation on code-switching text. While our models represent SoTA, the evaluation results indicate that it is a challenging task.
@@
Lyrics. Table 6 shows results on verse segmentation (i.e., segmenting songs into verse, chorus, etc.). None of the other baseline systems, including LLMs, can improve over the current domain-specific SotA, $\text{SSM}_{\text{string}}$. In contrast, SaT+LoRA outperforms $\text{SSM}_{\text{string}}$ by $10\%$ avg. F1. The difference is especially pronounced in hard-to-segment songs that are low in repetitiveness (e.g., Rap music), with a $15\%$ difference in F1 scores. When evaluating SaT+LoRA on manually corrupted lyrics, it still outperforms all baselines, even when compared to baselines evaluated on non-corrupted songs. Additionally, SaT+LoRA @1000, using 1000 songs per genre for adaptation, still outperforms all baselines. We provide complete results, including those for each genre, in Appendix A.4.
Legal. We provide comprehensive results on MultiLegalSBD in Appendix A.4.
@@ 7 Discussion
LLMs. Contrary to our expectations, our evaluation results reveal that LLMs generally underperform, particularly in non-English languages. Notably, when using LLMs for sentence segmentation via prompting, each sentence is processed twice – once as part of the input, appended to the prompt, and once within the output. This redundancy leads to inefficient processing, needing to copy the input verbatim to the output, ideally only adding newlines. However, in practice, LLMs are highly prone to alter their input (Barbero et al., 2024). We found this issue to be particularly severe for noisy text and lyrics. [^10] Llama 38B and Command R altered $1.5\%$ and $2\%$ of all characters within lyrics, respectively, even though we prompted them not to alter their input (c.f. Appendix A.2). This is highly problematic for a specific task requiring input and output characters to remain the same. Still, we tried to address this by using the Needleman-Wunsch sequence alignment algorithm to make pure segmentation performance comparable to other methods.
@@ 8 Conclusion
We proposed SaT, an efficient, robust, and highly adaptable multilingual sentence segmentation method that neither relies on language codes nor punctuation. Further, we introduced SaT+SM, improving SaT via supervised adaptation using multiple corruption schemes. Our method consistently achieves state-of-the-art performance in experiments across 85 languages and 8 diverse corpora, even outperforming newly introduced and optimized strong LLM baselines. We also demonstrated that SaT can be efficiently domain-adapted via LoRA, setting new performance standards on segmentation of lyrics and code-switching text. Overall, we hope SaT will unlock significantly improved text data (pre-)processing across a range of NLP applications for multiple languages and domains via its robust and consistently strong performance, versatility, and high efficiency.
@@ Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF): P36413, P33526, and DFH-23, and by the State of Upper Austria and the Federal Ministry of Education, Science, and Research, through grants LIT-2021-YOU-215. In addition, Ivan Vulić and Benjamin Minixhofer have been supported through the Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137) awarded to Ivan Vulić. This research has also been supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). This work was also supported by compute credits from a Cohere For AI Research Grant, these grants are designed to support academic partners conducting research with the goal of releasing scientific artifacts and data for good projects. We also thank Simone Teufel for fruitful discussions.
@@ A.2 Complete Experiment Details
For each language, we corrupt the data in two ways. In the first, we lowercase and remove all punctuation tokens. This aims to roughly emulate the output of an automatic speech recognition (ASR) system. In the second, we lowercase all text with probability $0.5$, remove all punctuation with probability $0.5$, and duplicate punctuation (e.g., changing ! to !!!) with geometric distribution scaling with the number of duplications (i.e., doubling with probability $0.5$, tripling with probability $0.25$, etc.). This aims to emulate user-generated text.
@@ A.4 Additional Results
For both SaT and WtP, we use a default stride of $64$ during evaluation. Since SaT operates on subwords but WtP on characters, this results in different scaling behaviors, illustrated in Figure 7.
@@
We show results using both language codes on code-switched text for models relying on language codes in Table 14.

### v2

@@ Abstract
Segmenting text into sentences plays an early and crucial role in many NLP systems. This is commonly achieved by using rule-based or statistical methods relying on lexical features such as punctuation. Although some recent works no longer exclusively rely on punctuation, we find that no prior method achieves all of (i) robustness to missing punctuation, (ii) effective adaptability to new domains, and (iii) high efficiency. We introduce a new model — Segment any Text (SaT) — to solve this problem. To enhance robustness, we propose a new pretraining scheme that ensures less reliance on punctuation. To address adaptability, we introduce an extra stage of parameter-efficient fine-tuning, establishing state-of-the-art performance in distinct domains such as verses from lyrics and legal documents. Along the way, we introduce architectural modifications that result in a threefold gain in speed over the previous state of the art and solve spurious reliance on context far in the future. Finally, we introduce a variant of our model with fine-tuning on a diverse, multilingual mixture of sentence-segmented data, acting as a drop-in replacement and enhancement for existing segmentation tools. Overall, our contributions provide a universal approach for segmenting any text. Our method outperforms all baselines — including strong LLMs — across 8 corpora spanning diverse domains and languages, especially in practically relevant situations where text is poorly formatted. [^1] Our models and code, including documentation, are available at https://github.com/segment-any-text/wtpsplit under the MIT license.
@@ 2 Introduction
However, the performance of WtP as the current state-of-the-art (SoTA) model and all other segmenters is still poor on texts from more challenging domains. This includes, among others, user-generated text such as tweets and highly heterogeneous domains such as lyrics. Segmenting these texts is challenging because of missing and/or extra punctuation, inconsistent spacing, and especially irregular casing. Furthermore, nearly all existing systems, including WtP, require the specification of the texts’ language at inference time. This necessitates an additional preprocessing step of language identification, which often proves to be imperfect, particularly with user-generated content (Lui and Baldwin, 2014; Sterner and Teufel, 2023). Moreover, this necessity limits their applicability to code-switching text.
@@
Figure 1 shows that the standard 3-layer version of SaT outperforms the current open weights state-of-the-art, WtP, while achieving a $\approx 3$x reduction in inference time. Overall, we present several innovations that overcome each of the shortcomings of previous methods, culminating in a universal model for sentence segmentation. We provide some examples of our model’s predictions in Figure 2.
@@
1) We introduce Segment any Text (SaT), an efficient method for sentence segmentation that can reliably segment text across 85 languages regardless of lexical features such as punctuation or casing. 2) We show how our models can be adapted to different domains via data-efficient means, requiring only a minimal set (e.g., 16) of sentence-segmented examples. 3) We train and release SaT models in five sizes, covering 85 languages, and demonstrate state-of-the-art performance across 8 corpora, even outperforming newly introduced strong (open weights) LLM baselines.
@@ 3.2 Where’s the Point (WtP)
Robustness to corruptions. To make WtP less reliant on punctuation, Minixhofer et al. (2023) randomly remove some punctuation during training. In addition, they predict the likelihood of commonly occurring punctuation as an auxiliary objective. For details, we refer to Appendix A.5. While this helps make WtP models less reliant on punctuation, we still find that WtP models have major issues when text is inconsistently formatted, especially irregular casing.
@@ 4 SaT: Segment any Text
Robustness to corruptions. We randomly remove common punctuation-only tokens with probability $p$ and use the auxiliary punctuation-prediction objective during training. For details, see § 3.2 and Appendix A.5. Further, we randomly remove all casing and punctuation in 10% of samples within a batch during training. The resulting model, Segment any Text (SaT), already shows strong segmentation performance at improved efficiency.
Still, to further improve SaT, we continue training it on a Supervised Mixture of already-segmented sentences. To be even less dependent on patterns such as punctuation, spaces, and casing, we augment the data by introducing several additional corruption schemes, resulting in our more specialized model, SaT+SM.
Our first corruption scheme removes all casing, if available, and punctuation tokens for all text. Secondly, we add randomness to the corruption in as many situations as we find useful, aiming to emulate user-generated text in tweets or forums. This includes duplicating punctuation, removing punctuation, lowercasing, and removing/adding spaces between sentences. Finally, we also use clean, non-corrupted text. We then sample uniformly across these three categories. For details, see Appendix A.2.
@@
Code- switching C.f. Table 10. Reddit posts for German-English (de-en); data for 3 additional language pairs taken by concatenating code-switching sentences from bilingual transcriptions. We treat all data as transcriptions, removing all punctuation and casing; we only keep sentences with at least one token of each language. Deuchar (2009), Osmelak and Wintner (2023), Nguyen and Bryant (2020) and Çetinoğlu (2017)
@@ 6.1 Performance on Clean Text
Table 2 shows evaluation results on clean text, averaged over OPUS100, UD, and Ersatz on a diverse selection of languages, including an average over 81 languages.666 For the average, we only consider languages with datasets with both train and test sets for a fair comparison. While we evaluate on 85 languages, this is the case in 81 languages. We categorize methods into (i) multilingual, which take only text as input, and (ii) monolingual, which additionally rely on a language code or, in the case of WtPT, WtPPUNCT, and SaT+LoRA, are adapted to a target domain.
@@ 6.3 Performance on Challenging Domains
Code-switching. The results in Table 5 reveal that WtP achieves an average F1 score of $29.1\%$, while the highest-performing LLM scores $43.4\%$. SaT and SaT+SM achieve average F1 of $48.2\%$ and $54.4\%$, respectively. SaT+LoRA continues to improve performance, achieving $61.7\%$. To the best of our knowledge, this is the first comprehensive evaluation of sentence segmentation tools on code-switching text. While our models now represent SoTA, the evaluation results indicate that it is a challenging task.
@@
Lyrics. Table 6 shows results on verse segmentation (i.e., segmenting songs into verse, chorus, bridge, etc.). None of the other baseline systems, including LLMs, can improve over the current domain-specific SotA, $\text{SSM}_{\text{string}}$. In contrast, SaT+LoRA outperforms $\text{SSM}_{\text{string}}$ by $10\%$ avg. F1. The difference is especially pronounced in hard-to-segment songs that are low in repetitiveness (e.g., Rap music), with a $15\%$ difference in F1 scores. When evaluating SaT+LoRA on manually corrupted lyrics, it still outperforms all baselines, even when compared to baselines evaluated on non-corrupted songs. Additionally, SaT+LoRA @1000, using 1000 songs per genre for adaptation, still outperforms all baselines. We provide complete results, including those for each genre, in Appendix A.4.
Legal and qualitative examples.. We provide comprehensive results on MultiLegalSBD in Appendix A.4. Finally, We provide qualitative examples from several domains in Appendix A.3.
@@ 7 Discussion
LLMs. Contrary to our expectations, our evaluation results reveal that LLMs generally underperform, particularly in non-English languages. Notably, when using LLMs for sentence segmentation via prompting, each sentence is processed twice – once as part of the input, appended to the prompt, and once within the output. This redundancy leads to inefficient processing, needing to copy the input verbatim to the output, ideally only adding newlines. However, in practice, LLMs are highly prone to alter their input (Barbero et al., 2024). We found this issue to be particularly severe for noisy text and lyrics. [^9] Llama 38B and Command R altered $1.5\%$ and $2\%$ of all characters within lyrics, respectively, even though we prompted them not to alter their input (c.f. Appendix A.2). This is highly problematic for a specific task requiring input and output characters to remain the same. Still, we tried to address this by using the Needleman-Wunsch sequence alignment algorithm to make pure segmentation performance comparable to other methods. [^10] The same objective could be achieved via other means, e.g., constrained decoding (Beurer-Kellner et al., 2024).
@@ 8 Conclusion
We proposed SaT, an efficient, robust, and highly adaptable multilingual sentence segmentation method that neither relies on language codes nor punctuation. Further, we introduced SaT+SM, improving SaT via supervised adaptation using multiple corruption schemes. Our method consistently achieves state-of-the-art performance among open weights models in experiments across 85 languages and eight diverse corpora, even outperforming newly introduced and optimized strong LLM baselines. We also demonstrated that SaT can be efficiently domain-adapted via LoRA, setting new performance standards on segmentation of lyrics and code-switching text. Overall, we hope SaT will unlock significantly improved text data (pre-)processing across a range of NLP applications for multiple languages and domains via its robust and consistently strong performance, versatility, and high efficiency.
@@ Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF): https://doi.org/10.55776/P33526, https://doi.org/10.55776/DFH23, https://doi.org/10.55776/COE12, https://doi.org/10.55776/P36413. In addition, Ivan Vulić and Benjamin Minixhofer have been supported through the Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137) awarded to Ivan Vulić. This research has also been supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). This work was also supported by compute credits from a Cohere For AI Research Grant, these grants are designed to support academic partners conducting research with the goal of releasing scientific artifacts and data for good projects. We also thank Simone Teufel for fruitful discussions.
@@ A.2 Complete Experiment Details
For each language, we corrupt the data in two ways. In the first, we lowercase and remove all punctuation tokens. This aims to roughly emulate the output of an automatic speech recognition (ASR) system. In the second, we lowercase all text with probability $0.5$, remove all punctuation with probability $0.5$, duplicate punctuation (e.g., changing ! to !!!) with geometric distribution scaling with the number of duplications (i.e., doubling with probability $0.5$, tripling with probability $0.25$, etc.), and join sentences without a whitespace with probability $0.1$ (or with a space for the four languages which do not generally use a whitespace to split sentences, see Table 1.) This aims to emulate user-generated text.
@@ A.4 Additional Results
For both SaT and WtP, we use a default stride of $64$ during evaluation. Each subword or character is thus processed multiple times, where we average predictions for overlapping positions. Since SaT operates on subwords but WtP on characters, this results in different scaling behaviors, illustrated in Figure 7.
@@
We show results using both language codes on code-switched text for models using language codes in Table 14.
@@
A.5 Auxiliary Punctuation Prediction Objective
As mentioned in Section 4, we adopt the auxiliary punctuation prediction objective from WtP (Minixhofer et al., 2023) as our base corruption scheme.
For clarity, we first specify the target without the auxiliary objective. Here, the original sequence of tokens $\bm{c}$ within some corpus is first stripped of newline characters:
[…]
We then create labels, which we set positive if the following token in the original sequence is a newline character:
$\bm{y}=\left\{\!\begin{array}[]{l}\!1\text{\; if $c_{i+1}=\mathtt{\backslash n}$}\\ \!0\text{\; otherwise}\end{array}\>|\>c_{i}\in\bm{x}\right\},$ (2)
where $c_{i}$ indexes into the original sequence $\bm{c}$. Using these labels, we optimize the standard cross-entropy of these labels and the model’s predictions. Note that the newline character is not contained in our base model’s vocabulary and will thus only appear as a single character. We also tokenize the whole batch at the start before applying any corruptions and do not re-tokenize later. We found this to be more effective than re-tokenizing the sequence after applying corruptions.
Auxiliary Punctuation Prediction. For the auxiliary objective, we adapt the methodology from Minixhofer et al. (2023) to tokens and identify the union of the 30 most common punctuation-only tokens within the training set. For simplicity, we ignore tokens containing multiple (potentially non-punctuation) characters. We also include the <UNK> token in this resulting set $P$, resulting in 109 punctuation tokens. We then define a random binary mask that determines which punctuation characters to remove among $P$, resulting in the new sequence $\bm{x}^{\prime}$:
$\bm{x}^{\prime}=\left\{c_{i}\>|\>\begin{array}[]{l}c_{i}\in\bm{c},c_{i}\neq\mathtt{\backslash n},\\ c_{i}\notin P\text{ or }p_{i}=0\\ \end{array}\right\}$ (3)
Here, we do not remove two consecutive character tokens to be able to reconstruct the original sequence. In addition, unlike WtP, we only remove character tokens if the following token is not a newline token. For the remaining characters, the auxiliary labels $\bm{z}$ indicate which (if any) character among $P$ followed them in the original sequence:
$\bm{z}=\left\{\!\begin{array}[]{l}\!c_{i+1}\text{\; if $c_{i+1}\in P$}\\ \!0\text{\;\;\;\;\;\;otherwise}\end{array}\>|\>c_{i}\in\bm{x}^{\prime}\right\}$ (4)
To avoid needing two separate forward passes through the model, we substitute the input $\bm{x}$ with $\bm{x}^{\prime}$ also for the main (newline prediction) objective. The final loss $\mathcal{L}$ is obtained by summing up the main newline prediction objective and the auxiliary objective of predicting punctuation:
$\mathcal{L}=\mathcal{L}^{\mathrm{main}}+\mathcal{L}^{\mathrm{aux}}$ (5)